Ernest is 4 times as old as Jacob. In 5 years' time ernest will be only 3 times as old as Jacob. If Jacob's age is x, form an equation in x and solve it to find how old ernest will be in 5 years' time.i
Answer by pj66(7)

Let x be Jacob's age.
So, Ernest's age is 4x
After 5 years,
Jacob's age is (x+5)
Ernest's age is 4x+5
=> 4x+5=3*(x+5)
=> 4x+5=3x+15
=> x=10
Ernest's age will be 45 in 5 years of time.